OpenStudy (anonymous):

Which expression is equivalent to 6a + 12? 6(a + 6) 3(3a + 9) 6(a + 9) 2(3a + 6)

pooja195 (pooja195):

Find a GCF for 12 and 6

pooja195 (pooja195):

Need help? :)

OpenStudy (anonymous):

yes

pooja195 (pooja195):

List the numbers the go into 6 and 12 obviously the GCF will be even since we are working with 2 even numbers :-) 2,4,6, <---our GCF GCF=6 :)

pooja195 (pooja195):

Now just divide the whole expression by 6 6/6=? 12/6=?

jimthompson5910 (jim_thompson5910):

you can also go through the options and check each one example: for choice A, we have 6(a + 6) distribute the outer 6 through to get 6(a + 6) = 6*a + 6*6 = 6a + 36

jimthompson5910 (jim_thompson5910):

6a + 36 is not the same as 6a + 12, so choice A is ruled out. Check the others the same way.